LHC Status Wed Morning 13-April R. Assmann, B. Holzer et al

Slides:



Advertisements
Similar presentations
8:16 SB 25ns dumped by RF; integrated lumi 0.6 nb-1. 9:14 BIC problem in TI8 and CMS recovering their tracker 10:09 Abort gap cleaning commissioning. 16:29.
Advertisements

Mo morning set up 228 bunch injection (36 bunch trains), ramp, squeeze, collisions h luminosity (depending on experiments) ATLAS length scale calibration.
Sunday :00 injection 336 bunches, ramp, squeeze luminosity. 03:19 Stable beams #1718. Duration: 9h08.  Peak lumi ATLAS/CMS: 3.8/3.6e32.  Int.
21/05/2012LHC 8:30 meeting Overview Week 20 Injectors, LHCb polarity, Beam-Beam Coordinators: Bernhard Holzer, Jan Uythoven.
LHC progress Report LHCC Wednesday 23 rd March 2011.
LHC Progress Friday 30 th October 2015 Coordination Week 44: Massimo Giovannozzi, Wolfgang Hofle, Jorg Wenninger.
Monday h34: beam (1020b) lost after 10 Hz RF trim. Fill h56: XPOC to be reset by expert. 08h18: Masked QPS_STATE SIS interlocks for RB.A12.
09:25 Operator dump fill #2178, integrated lumi 107/108/18 nb-1. Trip of RQ4.R6 (fire heaters discharged) just after the beam dump but the QPS gave no.
LHC Fri 6 & Sat 7 Nov :00 End instabilities MD, dump to refresh air at dump TDE cavern 08:00 Access for vacuum team at the TDE starting 10:30 Access.
Saturday 11.9 ● From Friday – Minimum required crossing angle is 100  rad in 2010 – Plenty of aperture at triplets: > 13  (n1 > 10) – Can stay with 170.
04h36: Cryo problem in arc 81. Access. 09h14: Injection.  TI8 injection collimator setup.  Testing new sequence.  New version of vacuum application.
07:00 Dump fil #2219, 123 pb -1 delivered. Trim TDI parking position to +/- 55 mm in the collimator BP trimmed the temperature kicker limit to 62 degrees.
Monday 19 th March 07:33 Lost (another) beam in the squeeze, beyond 2 m, B2 hits 1/3 resonance. 08:30 Start (another) ramp  Try to correct coupling and.
LHC Status Sat Morning 30-June Bernhard Holzer, Jan Uythoven Status Friday Late: preparing restart some issues... LHCb dipole tripped / router problem.
:30 meetingEBH 1  No beam during the morning:  4:15 – 9: V distributor: Recovery after electrical interruption  access for TI services.
LHC Status Fri Morning 3-June
Ralph Assmann, Giulia Papotti, Frank Zimmermann 25 August 2011
MD Planning Fri – Sat (26. – 27.8.)
Saturday Coordinators weekend: Mike Lamont, Eva Barbara Holzer
Wednesday :06 Dump fill 3533: interlock BPMS in IP6
Cryo Problem MD Planning Tue (1.11.) C B Day Time MD MP Tue 01:00
Wednesday September 14th
Monday 14th March Second fill with pilot + 3 nominal bunches
Monday 11/07: Recovery Plenty of accesses.
Week 46 Week 46: Machine coordinators: Roger Bailey – Gianluigi Arduini Main aims of the week: Stable beams with ions Scheduled stop for ion source refill.
Tue :25 Beams dumped due to RF vacuum interlock
LHC Status Su Morning 17-April R. Assmann, B. Holzer et al
Summary of week 44 Aims of week 44 J. Wenninger & J. Uythoven
LHC Morning Meeting - G. Arduini
Summary of Week 13 G. Arduini, B. Holzer
Summary of Week 16 G. Arduini, J. Wenninger
LHC Morning Meeting - G. Arduini
Thursday 10th May - morning
Saturday 2/4/ :00-09:00 Problems with SPS:
Wednesday 8th August & night
Friday March 4th 08h32: Access for beam dump system.
Friday 14th December - morning
MD#2 News & Plan Tue – Wed (19. – 20.6.)
Tuesday 28th September Fill 1375 ongoing 09:00 Move in Roman Pots
Wire almost entirely through beam before BLM triggers dump
Finally.... We had to wait a long time.... We had many problems....
Friday Morning 7:30 injection for fill 2908
Saturday 7th May Sat – Sun night
Summary Thursday h21: Stable beams fill #1303.
Wednesday 10/5 Cryo recovery until 18:30
Sunday/Monday & :30 Stable beams. Fill #1851.
Wednesday Ramp single nominal bunches for loss maps
Aims for the week Recover from technical stop and precycle at 10A/s
Monday 14:00 beam dump Access:
Monday 8th December Morning used for RF work (IP z, injection), damper noise checks. Stable beams period with LHCb spectrometer OFF – lost beam after.
Friday 8/10/10 - Morning Planned to keep fill until the afternoon
Monday morning 09:40 Dump fill 2838, integrated ~130 pb-1.
Summary for LPC (March 7th, Ralph Assmann)
Saturday :42 stable beams fill 3541
Saturday 12 March Morning: problem on RSD1.A12B2 – flow rate problem on the cooling water for the converter. 14:00: Loss map -500 Hz (positive energy offset)
LHC Morning Meeting - G. Arduini
Tuesday 20 September 2011.
Wednesday 23/2 Thanks CRYO!!!.
Wednesday 6th April Injection of 50 ns up to 108 bunches (3 x 36).
Saturday 24 September 2011.
Mon 2/4 08:00-13:00: Access for: Experiments Power converters BBQ
Wednesday h18: 228b fill lost just before start of ramp.
Monday 26th September 7:30 Beam dump, QSP on RQX.R1, lost cryo in ITR1, 91 pb-1. Cryo recovery, needed about 1.5 h. Setting of TI2 collimators to 5 sigma.
Friday 15th February - morning
Another Immortal Fill….
Monday 10th December - morning
Wed 12/10 09:30 – 10:30 Ramp down of ALICE solenoid to 50% nominal field on request of ALICE. Increase of the vacuum (e-cloud in this area that has not.
Monday 152 bunch operation summary
Saturday Morning used to complete the loss maps and asynch dumps (with RPs). All loss maps OK (B. Salvachua), but the off-p cleaning for B1 (Lead) was.
Tuesday
Presentation transcript:

LHC Status Wed Morning 13-April R. Assmann, B. Holzer et al

Tentative Plan for Monday / Tuesday ✔ ?

Tue Morning: Restart after tech. Problems Main techniccal issues: Access system: e-stopper relais Fire detection system Access door got a new dump due to an access door in P4 which opened by itself. QPS quench heaters Restart: injection setup 28/03/2011 LHC 8:30 meeting

Tue Morning: e-cloud checks RF frequency trimmed by +10Hz +/- 10 Hz but ... beams dumped - high losses in IP7 28/03/2011 LHC 8:30 meeting

Tue Morning: the 10 Hz problem e-cloud ? IR 7 beams dumped - high losses in IP7 LSS 5 / 1 28/03/2011 LHC 8:30 meeting

Tue Morning: the 10 Hz problem loss pattern: b2 Loss occuring ~5 minutes after the frequency trim The lossess affect initially B2 then B1 again and finally B2 The first loss is affecting all bunches while the second only some of the bunches of the tail of the trains No sign of instability is observed either in the BBQ nor from the damper pick-ups Losses at TCPs are consistent with an increase in abort gap population 28/03/2011 LHC 8:30 meeting

Tue Morning: the 10 Hz problem en detail: b2 beam current frequency abort gap intensity bbq amplitude y bbq amplitude y 28/03/2011 LHC 8:30 meeting

Tue Morning: the 10 Hz problem en detail: b1 beam current frequency abort gap intensity bbq amplitude y bbq amplitude y 28/03/2011 LHC 8:30 meeting

Tue Morning: the 10 Hz problem en detail: b1 beam current b1 frequency beam current b2 28/03/2011 LHC 8:30 meeting

Tue Morning: the 10 Hz problem da capo: 48 bunches in... rimmed again + 10 Hz. Nothing happening. No additional losses. 228 bunches circulating. Switching off abort gap cleaning. Trimmed +10 Hz. Nothing happening. Abort gap population went up when reverting frequency trim. Sent another 10 Hz (total + 20 Hz off). Another step in the abort gap population. 28/03/2011 LHC 8:30 meeting

Tue Late: Restart after tech. Problems And again: technical issues: BLM intervention status: Identified high voltage supply fault on the whole arc of R4 21:30h Ventilation door interlock for injection. -> Access ... Access finished, door had been really open. 0:31-3:30h Piquet working on PSB LLRF. 28/03/2011 LHC 8:30 meeting

Tue Night: scrubbing checks 1.) Restart after technical Problems inject 2 nominal & 1 pilot per beam precycle, inject 36 bunches. ramp & squeeze: test the + /- 10 Hz problem bunches ? how much space do we have in the basic beam parameters tune, orbit, chroma, coupling etc Trimmed QV & QH +0.003 for both beams. Trimmed RF by -20 Hz (at nominal tunes).

Wed Morning: 2.) Test Ramp to 3.5 TeV: 48 bunches: inject and ramp 48 bunches in both beams, ramp, squeeze, collisions ... observe vacuum, rf, cryo load etc check for beam robustness:apply + / - 20 Hz change tune by + / - 0.003 in both beams scan 3.) 200 bunches: inject ,ramp, squeeze, collisions ... observe vacuum, rf, cryo load etc check for beam robustness: 1 hour for the experiments 4.) rescrubbing, inject 1020 bunches ,go on scrubbing at 450 GeV observe vacuum & cryo load ✔ ✔